# Temporally Coherent Imitation Learning via Latent Action Flow Matching for Robotic Manipulation

### v1

@@ IV-C1 Real-World Experimental Setup
Expert demonstrations are collected via human teleoperation, with 30 demonstrations per task. We compare LG-Flow Policy against DP3 [32], Reactive Diffusion Policy (RDP) [28], and Flow Policy [34]. For RDP, tactile inputs are replaced with wrist camera images to ensure a fair comparison under visual-only settings.
@@ IV-C2 Real-World Results Analysis
LG-Flow Policy (Ours) $\mathbf{83.3\pm 5.8}$ $\mathbf{96.7\pm 5.8}$ $\mathbf{60.0\pm 0.0}$ $\mathbf{70.0\pm 5.8}$ $\mathbf{77.5}$ $8.59\text{ ms}$
@@
As shown in Fig. 6, LG-Flow Policy achieves the lowest mean and maximum smoothness scores among all methods, indicating substantially reduced jerk and high-frequency oscillations. Compared to Flow Policy, the mean smoothness metric is reduced by over $93.7\%$, demonstrating strong robustness under real-world sensing noise and actuation delays. Consistent with simulation results, RDP exhibits lower smoothness than other diffusion-based baselines, suggesting that latent action representations improve execution stability in physical settings. However, its trajectories remain less smooth than those of LG-Flow Policy, as RDP does not impose explicit temporal smoothness constraints in the latent space.
This effect is further illustrated in Fig. 7, where Flow Policy produces pronounced high-frequency oscillations, while LG-Flow Policy yields smooth and temporally coherent joint motions under identical initial conditions.
Figure 7: Comparison of real-world joint trajectories under identical initial conditions. Flow Policy exhibits high-frequency oscillations and abrupt variations, while LG-Flow Policy generates smoother and more temporally coherent trajectories, consistent with the quantitative smoothness metrics.
@@
Table II reports task success rates and inference latency in real-world experiments. LG-Flow Policy achieves the highest average success rate of $77.5\%$, outperforming DP3 ($57.9\%$), RDP ($70.0\%$), and Flow Policy ($52.5\%$). While Flow Policy attains the lowest latency (6.54 ms), its unstable trajectories often lead to execution failures. LG-Flow Policy incurs only a modest increase in response time (8.59 ms) while achieving substantially higher task success. One contributing factor to the improved success of both RDP and LG-Flow Policy is the incorporation of wrist-mounted camera observations during action decoding, providing execution-time visual feedback for precise manipulation. LG-Flow Policy remains over $3\times$ faster than DP3 and RDP, indicating that significant real-world performance gains are achieved at minimal computational cost.
Overall, real-world results confirm that LG-Flow Policy strikes a robust balance between inference efficiency, trajectory smoothness, and task success, supporting its suitability for stable real-world deployment.
@@ IV-D Ablation Study
Figure 8: Ablation study on trajectory smoothness and task success rate in real world manipulation tasks.
Fig. 8 summarizes ablation results in both simulation and real-world settings, analyzing the contributions of geometry-aware perception (GE-Flow) and latent action modeling (L-Flow). Across both domains, the two components exhibit complementary effects on task success and trajectory smoothness.
In simulation, geometry-aware encoding improves average success from $61\%$ to $64.3\%$ but slightly degrades trajectory smoothness, indicating that enhanced spatial perception alone does not resolve temporal instability. In contrast, latent action modeling substantially reduces the smoothness metric ($0.228\rightarrow 0.046$) while increasing success to $72.4\%$. Combining both components further improves success to $78.3\%$ while maintaining low oscillations.
A similar trend is observed in real-world experiments. While geometry-aware perception alone yields only marginal success gains ($52.5\%\rightarrow 54.2\%$), it noticeably increases trajectory oscillations, often leading to execution failures under real-world noise and actuation delays. Latent action modeling, by contrast, improves both metrics, increasing success to $71.7\%$ and reducing smoothness by over $95\%$. Combining both components achieves the best real-world performance, reaching a success rate of $77.5\%$ with minimal oscillations.
Overall, these results indicate that geometry-aware perception enhances spatial understanding, whereas stable manipulation critically relies on enforcing temporal coherence through latent action modeling, particularly in real-world settings.
@@ V Conclusion
This work addresses real-time robotic manipulation under long-horizon and high-dimensional control, where existing generative policies struggle to jointly achieve expressive behavior modeling, fast inference, and stable execution, limiting their deployment on physical systems.
We propose LG-Flow Policy, a trajectory-level action generation framework that performs flow matching in a continuous latent action space. By mapping action sequences into a temporally regularized latent trajectory manifold and learning an explicit latent-space flow, LG-Flow Policy enables near single-step inference while substantially improving execution smoothness and stability. Operating on motion-level latent representations decouples global trajectory structure from low-level control noise, mitigating error amplification during generation and leading to more reliable long-horizon behaviors.
LG-Flow Policy further integrates geometry-aware point cloud conditioning and execution-time decoding modulation, allowing additional sensory signals to be incorporated without perturbing latent trajectory dynamics. This design enables richer multimodal execution while preserving fast response, as demonstrated by visual-conditioned decoding with a wrist-mounted camera in real-world experiments. Extensive simulation and real-robot results show smooth trajectories, high success rates, and rapid inference, outperforming both diffusion-based and raw action-space flow-based baselines.
Despite its effectiveness, the current system adopts relatively lightweight multimodal fusion and does not explicitly model complex contact dynamics. Future work will explore richer multimodal representations and contact-aware latent dynamics to extend LG-Flow Policy toward more challenging contact-rich manipulation.

### v2

@@ IV-C1 Real-World Experimental Setup
Expert demonstrations are collected via human teleoperation, with 30 demonstrations per task. We compare CoLA-Flow Policy against DP3 [33], Reactive Diffusion Policy (RDP) [28], and Flow Policy [35]. For RDP, tactile inputs are replaced with wrist camera images to ensure a fair comparison under visual-only settings.
@@ IV-C2 Real-World Results Analysis
CoLA-Flow Policy $\mathbf{83.3\pm 5.8}$ $\mathbf{96.7\pm 5.8}$ $\mathbf{60.0\pm 0.0}$ $\mathbf{70.0\pm 5.8}$ $\mathbf{77.5}$ $8.59\text{ ms}$
@@
As shown in Fig. 6, CoLA-Flow Policy achieves the lowest mean and maximum smoothness scores, indicating substantially reduced jerk and high-frequency oscillations in real-world execution. Compared to Flow Policy, the mean smoothness is reduced by over $93.7\%$, demonstrating robustness to sensing noise and actuation delays. Consistent with simulation results, RDP improves smoothness over other diffusion-based baselines, but remains less smooth than CoLA-Flow Policy due to the lack of explicit temporal smoothness constraints in the latent space.
This effect is further illustrated in Fig. 7, which shows joint trajectories over time under identical initial conditions. Flow Policy exhibits pronounced high-frequency oscillations, whereas CoLA-Flow Policy yields smoother, temporally coherent joint motions.
Figure 7: Comparison of real-world joint trajectories under identical initial conditions. Flow Policy exhibits high-frequency oscillations and abrupt variations, while CoLA-Flow Policy generates smoother and more temporally coherent trajectories, consistent with the quantitative smoothness metrics.
@@
Table II reports task success rates and inference latency in real-world experiments. CoLA-Flow Policy achieves the highest average success rate of $77.5\%$, outperforming DP3 ($57.9\%$), RDP ($70.0\%$), and Flow Policy ($52.5\%$). While Flow Policy attains the lowest latency (6.54 ms), its unstable trajectories often lead to execution failures. CoLA-Flow Policy incurs only a modest increase in response time (8.59 ms) while achieving substantially higher task success. One contributing factor to the improved success of both RDP and CoLA-Flow Policy is the incorporation of wrist-mounted camera observations during action decoding, providing execution-time visual feedback for precise manipulation. CoLA-Flow Policy remains over $3\times$ faster than DP3 and RDP, indicating that significant real-world performance gains are achieved at minimal computational cost. Overall, real-world results confirm that CoLA-Flow Policy strikes a robust balance between inference efficiency, trajectory smoothness, and task success, supporting its suitability for stable real-world deployment.
@@ IV-D Ablation Study
Figure 8: Ablation study on trajectory smoothness and task success rate in real-world manipulation tasks.
Fig. 8 summarizes ablation results in both simulation and real-world settings, analyzing the contributions of geometry-aware perception (GE-Flow) and latent action modelling (L-Flow). Across both domains, the two components exhibit complementary effects on task success and trajectory smoothness.
In simulation, geometry-aware encoding improves average success from $61\%$ to $64.3\%$ but slightly degrades trajectory smoothness, indicating that enhanced spatial perception alone does not resolve temporal instability. In contrast, latent action modelling substantially reduces the smoothness metric ($0.228\rightarrow 0.046$) while increasing success to $72.4\%$. Combining both components further improves success to $78.3\%$ while maintaining low oscillations.
A similar trend is observed in real-world experiments. While geometry-aware perception alone yields only marginal success gains ($52.5\%\rightarrow 54.2\%$), it noticeably increases trajectory oscillations, often leading to execution failures under real-world noise and actuation delays. Latent action modelling, by contrast, improves both metrics, increasing success to $71.7\%$ and reducing smoothness by over $95\%$. Combining both components achieves the best real-world performance, reaching a success rate of $77.5\%$ with minimal oscillations.
Overall, these results indicate that geometry-aware perception enhances spatial understanding, whereas stable manipulation critically relies on enforcing temporal coherence through latent action modelling, particularly in real-world settings.
@@ V CONCLUSION
This work addresses real-time robotic manipulation under long-horizon and high-dimensional control, where existing generative policies struggle to jointly achieve expressive behavior modelling, fast inference, and stable execution, limiting their deployment on physical systems.
We propose CoLA-Flow Policy, a trajectory-level action generation framework that performs flow matching in a coherent latent action space. By mapping action sequences into a temporally coherent latent trajectory manifold and learning an explicit latent-space flow, CoLA-Flow Policy enables near single-step inference while substantially improving execution smoothness and stability. Operating on motion-level latent representations decouples global trajectory structure from low-level control noise, mitigating error amplification during generation and leading to more reliable long-horizon behaviours.
CoLA-Flow Policy further integrates geometry-aware point cloud conditioning and execution-time decoding modulation, allowing additional sensory signals to be incorporated without perturbing latent trajectory dynamics. This design enables richer multimodal execution while preserving fast response, as demonstrated by visual-conditioned decoding with a wrist-mounted camera in real-world experiments. Extensive simulation and real-robot results show smooth trajectories, high success rates, and rapid inference, outperforming both diffusion-based and raw action-space flow-based baselines.
Despite its effectiveness, the current system adopts relatively lightweight multimodal fusion and does not explicitly model complex contact dynamics. Future work will explore richer multimodal representations and contact-aware latent dynamics to extend CoLA-Flow Policy toward more challenging contact-rich manipulation.